# Multi-mode Pulsations in AGB Stars: Insights from 3D RHD CO5BOLD Simulations

### v1

@@ 3.2 Radial modes and nodes
3.3 Mean density relation
Figure 9: Comparison of derived pulsation periods for different degrees; which is also referred to as a Petersen diagram. Top panel: Logarithmic ratio of periods derived from $\ell=2\,\mathrm{and}\,\ell=0$. Bottom panel: Logarithmic ratio of periods derived from $\ell=1\,\mathrm{and}\,\ell=0$. The colours indicate pulsation modes in the fundamental, first and second overtones, represented by red, green, and blue, respectively. The contours in each main panel illustrates the data point distribution. The subpanels in each row show histograms of the ratios, with dashed lines marking the mean values. For the bottom panel, the blue dashed line in the histogram represents the mean ratio when a luminosity criterion of $\log(\mathrm{L}_{\star}\,[\mathrm{L}_{\odot}])\leq 3.5$ is applied to the models being considered for the corresponding ratio results.
Figure 10: Relative power fractions of the radial, dipole, and quadrupole modes as a function of stellar mean density. The top row shows the relative power distribution across radial orders for each angular degree $\ell$, calculated as ${P^{\ell}_{n}}/{\sum_{\ell^{\prime}}P^{\ell^{\prime}}_{n}}$. The black-lines in the leftmost panel indicate the fractions of the growth rates of the fundamental, first and second overtones relative to summed-up growth rates, providing a means of comparing the relative amplitudes derived from linear 1D models (Trabucchi et al. 2019), as described in Sect. 3.1.1. The bottom row displays the relative power for each angular degree at a given radial order $n$, calculated as ${P^{\ell}_{n}}/{\sum{n^{\prime}}P^{\ell}{n^{\prime}}}$. These plots illustrate how the dominant mode shifts depending on the mean density and radial order.
@@ 3.4 Comparing radial and non-radial modes
3.4.1 Period hierachy of modes
@@ 3.5 Wavelets comparison between observations and models
Our analysis here focuses on mode-switching behaviour and pulsation periods, as a detailed investigation of pulsation modulation and power is beyond the scope of this work. The WWZ transform was applied to the visual light curves of R Dor, RY Dra, and AF Cyg, as acquired from the AAVSO database, with results shown in Fig. 11. To compare with the observations, we applied the WWZ transform to the time-series luminosity of the 3 representative models described in Sect. 3.2. Their time-series luminosity and respective WWZ spectra are displayed in Fig. 12.
@@ 4.2 Mode switching between radial and non-radial modes
As shown in Fig. 10, our grid of models reveals systematic trends in the dominant pulsation modes as a function of stellar parameters. In our framework (which follows definition of the sequences described in Trabucchi et al. 2017), sequence A corresponds to the second overtone, while the fundamental and first overtone modes are associated with sequences C and C’ (or B), respectively. Observational studies have reported differing results regarding which mode dominates at various luminosities and radial orders. For example, Wood (2015) and Stello et al. (2014) find that dipole ($\ell=1$) modes dominate sequence A, particularly at higher luminosities. This broadly aligns with Yu et al. (2020), who also report dipole dominance in overtone sequences. In contrast, Mosser et al. (2013) observe that radial ($\ell=0$) modes dominate at low pulsation frequencies ($\leq 1.0\,\mu\mathrm{Hz}$), corresponding to high-luminosity stars, particularly in the fundamental and first overtone sequences.
@@ 4.3 Amplitude and excitation of multi-mode pulsations
The top of the convection zone of AGB stars is covered by small-scale convection cells analogue to solar granules. However, in AGB stars, these cells are often obscured by cool and dense clouds of matter just above (Freytag et al. 2024). The downdrafts starting at the borders of the granules merge in layers below the surface and form tree-like structures building up even larger configurations in the deep interior. Some downdrafts reach from the surface to the core of the models. In this core region, low-order non-radial long-lived (several pulsation periods) convective flows exists. These connect to the downdrafts from above, which can experience an extra acceleration close to the core due to the drop in gravitational potential. The peak velocity in the downdrafts exceeds the typical upflow velocities by a margin and is larger than the RMS velocity (see Fig. 13). As a result, for the higher-luminosity – most Mira-like – models, the transport timescale within a downdraft from the surface to the model core is slightly shorter than the sound-crossing timescale. Velocities and Mach numbers of convective flows and radial fundamental acoustic pulsations are large, and their temporal and spatial scales are similar. Thus, the conditions for interactions are favourable.
@@
In a simple piston model for the excitation of acoustic pulsations via the $\kappa$ mechanism, the opacity is impacted by the periodically varying temperature and pressure. This modulates the radiative energy transport leading to temperature and pressure variations of the gas below the piston in excess to what is expected from adiabatic pulsations alone. In our 3D AGB models, however, radiative energy transport in the interior plays hardly any role. Therefore, possible modulation effects from the $\kappa$ mechanism would have no consequences. Still, there are other mechanisms that affect the pressure.
@@
A central pressure enhancement exactly in phase with the adiabatic pressure variation would lead to an enlarged restoring force and thus to a faster pulsation. A pressure increase slightly out of phase could either lead to a driving of the pulsations (if occurring in the expansion phase) or a damping (if occurring during the compression phase) of the mode. Convective fluxes vary during a pulsation cycle. During the compression phase, radial mass, momentum, and energy fluxes are strongly enlarged. They all affect the total central pressure. An enhanced mass flux could increase the gas pressure. The same happens when the (downward directed) flux of kinetic energy grows. In contrast, an enhanced (upward) enthalpy flux would decrease the central gas pressure. Additionally, transport of momentum could increase the central dynamical pressure. While all the contributions are roughly in phase with the pulsation cycle, the exact timings vary slightly – relative to the pulsation phase and with respect to each other, depending on the details of the convective flow structure. Short-term changes in the configuration of the large-scale convections cells can cause strong variations in the pulsations amplitude. On the other hand, a long-lasting convective dipole flow, for example, leads to stable, large-amplitude pulsations.
@@
Wavelet analyses of both our models and observations (see Sect. 3.5) reveal mode-switching behaviour, with overtone modes appearing less stable than the fundamental mode. As stars evolve up the AGB, the interaction of convection and pulsations changes significantly. Pulsations of higher-density, lower-luminosity AGB stars share similarities with solar-like oscillations, where turbulent pressure fluctuations in the outer convective envelope excite the pulsations. The situation differs in more luminous AGB stars. In these stars, deeper and larger-scale convective structures within a lower-density interior become prominent. Strong convective downdrafts can drive high-amplitude radial fundamental modes when in phase with pulsations but can also suppress multiple modes due to their disruptive nature (as discussed in the preceding paragraphs). Conversely, in stars with weaker convection, where motions do not reach as deeply, pressure maxima between the core and surface may help sustain overtone modes by favouring stable oscillatory patterns.
@@
To elucidate the differences in the timescales of the convective and acoustic pulsations, Fig. 13 shows the logarithmic ratio of convective velocity (computed as the RMS of radial velocities) to the local sound speed as a function of radius. The colour-coded profiles indicate that in the deep interior, where density is high, convective speeds are low, and sound speeds are high. Near the stellar surface, convective speeds increase while sound speeds decrease. Stars with higher mean densities consistently show lower convective velocities and higher sound speeds throughout their interiors. The results suggest that overtone amplitudes are more likely to be sustained in stars with extended inner regions where convective velocities are significantly lower than the local sound speed. Although it is difficult to determine a precise threshold for this ratio, the overall trend indicates that higher overtone amplitudes correlate with low convective velocities relative to sound speeds.
@@
The amplitude of multiple pulsation modes depends on a balance between excitation and damping processes. MacLeod et al. (2023) interpreted their results in terms of the survival of radial nodes, sustaining overtone modes. In stars with vigorous convection, such as AGB stars and red supergiants like Betelgeuse, high convective velocities may destabilise these nodes, contributing to mode-switching and irregular pulsation behaviour. This in part may help explain the observed mode-switching in R Dor and other red giants that exhibit irregular multi-mode pulsation behaviour.

### v2

@@ 3.3 Mean density relation
Figure 10: Comparison of derived pulsation periods for different degrees. Top panel: Logarithmic ratio of periods derived from $\ell=2\,\mathrm{and}\,\ell=0$. Bottom panel: Logarithmic ratio of periods derived from $\ell=1\,\mathrm{and}\,\ell=0$. The colours indicate pulsation modes in the fundamental, first and second overtones, represented by red, green, and blue, respectively. The contours in each main panel illustrates the data point distribution. The subpanels in each row show histograms of the ratios, with dashed lines marking the mean values. For the bottom panel, the blue dashed line in the histogram represents the mean ratio when a luminosity criterion of $\log(\mathrm{L}_{\star}\,[\mathrm{L}_{\odot}])\leq 3.5$ is applied to the models being considered for the corresponding ratio results.
Figure 11: Relative power fractions of the radial, dipole, and quadrupole modes as a function of stellar mean density. The top row shows the relative power distribution across radial orders for each angular degree $\ell$, calculated as ${P^{\ell}_{n}}/{\sum_{\ell^{\prime}}P^{\ell^{\prime}}_{n}}$. The black-lines in the leftmost panel indicate the fractions of the growth rates of the fundamental, first and second overtones relative to summed-up growth rates, providing a means of comparing the relative amplitudes derived from linear 1D models (Trabucchi et al. 2019), as described in Sect. 3.1.1. The bottom row displays the relative power for each angular degree at a given radial order $n$, calculated as ${P^{\ell}_{n}}/{\sum{n^{\prime}}P^{\ell}{n^{\prime}}}$. These plots illustrate how the dominant mode shifts depending on the mean density and radial order. Note that all power values are evaluated at the stellar surface for each identified mode.
3.3.1 Dominant pulsation modes
@@ 3.4 Comparing radial and non-radial modes
3.4.1 Period hierarchy of modes
@@ 3.5 Wavelets comparison between observations and models
Our analysis here focuses on mode-switching behaviour and pulsation periods, as a detailed investigation of pulsation modulation and power is beyond the scope of this work. The WWZ transform was applied to the visual light curves of R Dor, RY Dra, and AF Cyg, as acquired from the AAVSO database, with results shown in Fig. 12. To compare with the observations, we applied the WWZ transform to the time-series luminosity of the three representative models described in Sect. 2.2. Their time-series luminosity and respective WWZ spectra are displayed in Fig. 13.
@@ 4.2 Mode switching between radial and non-radial modes
As shown in Fig. 11, our grid of models reveals systematic trends in the dominant pulsation modes as a function of stellar parameters. In our framework (which follows definition of the sequences described in Trabucchi et al. 2017), the fundamental mode corresponds to sequence C, the first overtone spans sequences C’ and B, and the second overtone is associated with sequence A. Observational studies have reported differing results regarding which mode dominates at various luminosities and radial orders. For example, Wood (2015) and Stello et al. (2014) find that dipole ($\ell=1$) modes dominate sequence A, particularly at higher luminosities. This broadly aligns with Yu et al. (2020), who also report dipole dominance in overtone sequences. In contrast, Mosser et al. (2013) observe that radial ($\ell=0$) modes dominate at low pulsation frequencies ($\leq 1.0\,\mu\mathrm{Hz}$), corresponding to high-luminosity stars, particularly in the fundamental and first overtone sequences.
@@ 4.3.1 Convection and its pressure feedback
The surface of the convection-zone of AGB stars is covered by small-scale convection cells analogous to solar granules. However, in AGB stars, these cells are often hidden beneath cool, dense clouds of matter just above (Freytag et al. 2024). The downdrafts that originate along the granule borders merge below the surface and form tree-like structures that extend still further inward. Some downdrafts descend all the way from the surface to the model core. In this core region, low-order non-radial, long-lived (several pulsation periods) convective flows exist and connect with the downdrafts from above; the latter experience extra acceleration near the core because of the falling gravitational potential. The peak velocity in the downdrafts therefore exceeds both the typical up-flow velocity (see Fig. 3 in Ahmad et al. 2023) and the global RMS velocity. Figure 14 plots the ratio of the RMS convective speed to the local sound speed as a function of radius; if the same plot were computed with downdraft velocities, the entire curve would be shifted upward, reflecting the systematically higher downdraft velocities. Velocities and Mach numbers of convective flows and radial fundamental acoustic pulsations are large, and their temporal and spatial scales are similar. Thus, the conditions for interactions between convection and pulsations are favourable (Ahmad et al. 2023).
@@
In a simple piston model, where the excitation of acoustic pulsations is achieved via the $\kappa$-mechanism, the opacity is impacted by the periodically-varying temperature and pressure. This modulates the radiative energy transport leading to temperature and pressure variations of the gas below the piston in excess to what is expected from adiabatic pulsations alone. In our 3D AGB simulations, radiative transport inside the star is negligible, so any modulation predicted by considering $\kappa$-mechanism is inconsequential. Other mechanisms must therefore alter the pressure, evidenced by non-linear changes in pulsation amplitudes.
@@
A central pressure enhancement exactly in phase with the adiabatic pressure variation increases the restoring force and accelerates the pulsation (i.e. a faster pulsation). A pressure increase slightly out of phase could either lead to a driving of the pulsations (if occurring in the expansion phase) or a damping (if occurring during the compression phase) of the mode. Convective fluxes vary during a pulsation cycle. During the compression phase, radial mass, momentum, and energy fluxes are strongly enlarged. They all affect the total central pressure. An enhanced mass flux could increase the gas pressure. The same happens when the (downward directed) flux of kinetic energy grows. In contrast, an enhanced (upward) enthalpy flux would decrease the central gas pressure. Additionally, transport of momentum could increase the central dynamical pressure.
Although the contributions of pressure enhancements and convective fluxes vary roughly in phase with the pulsation, their exact timings differ, depending on the detailed convective geometry. Short-term changes in the configuration of the large-scale convections cells can cause strong variations in the pulsations amplitude. On the other hand, a long-lasting convective dipole flow, for example, leads to stable, large-amplitude pulsations.
4.3.2 Mode switching and quantitative diagnostics
@@
Wavelet analyses of both our models and observations (see Sect. 3.5) reveal mode-switching behaviour, with overtone modes appearing less stable than the fundamental mode. As stars evolve up the AGB, the interaction of convection and pulsations changes significantly. Pulsations of higher-density, lower-luminosity AGB stars share similarities with solar-like oscillations, where turbulent pressure fluctuations in the outer convective envelope excite the pulsations. The situation differs in more luminous AGB stars. In these stars, deeper and larger-scale convective structures within a lower-density interior become prominent. Strong convective downdrafts can drive high-amplitude radial fundamental modes when in phase with pulsations, but can also suppress overtone modes due to their disruptive nature (as discussed in Sect. 4.3.1). Conversely, in stars with weaker convection, where motions do not reach as deeply, pressure maxima between the core and surface may help sustain overtone modes by favouring stable oscillatory patterns.
@@
To elucidate the differences in the timescales of the convective and acoustic pulsations, Fig. 14 shows the logarithmic ratio of convective velocity (computed as the RMS of radial velocities) to the local sound speed as a function of radius. In the dense interior convective speeds are low and sound speeds high, whereas near the surface convective speeds rise and sound speeds fall. Stars of higher mean density therefore display lower convective velocities and higher sound speeds throughout. Altogether, Fig. 14 suggests that overtone amplitudes persist where convective speeds remain well below the local sound speed; however, it is difficult to determine a precise threshold for this.
@@
4.3.3 Sustaining radial nodes and an analogy with Cepheids
The amplitude of multiple pulsation modes depends on a balance between excitation and damping processes. MacLeod et al. (2023) suggested that the persistence of overtone modes in Betelgeuse depends on the survival of radial nodes, and that strong-enough convection may erase these nodes and trigger mode switching. By extending this idea, the high convective velocities sustained in AGB stars may destabilise any radial nodes; contributing to mode-switching and irregular pulsation behaviour. This in part may help explain the observed mode-switching in evolved red giants.